OpenStudy (unklerhaukus):

The square of any odd number is 1 more than a multiple of 8 true or false

OpenStudy (unklerhaukus):

is this case, can zero be considered as a multiple of eight ?, or is this the only exception , which would make the original statement false

OpenStudy (unklerhaukus):

@TuringTest

OpenStudy (jiteshmeghwal9):

\[5^2=25\]\[\implies (8 \times 3)+1=25\]so true.

OpenStudy (unklerhaukus):

\[1^2-1=0\times8\], but is zero a multiple of eight ?

OpenStudy (unklerhaukus):

i am assuming 1 is an odd number ,

OpenStudy (jiteshmeghwal9):

i think yes because after multiplying with zero the result is coming 0 so not surely but i think that 0 is the multiple of 8.

OpenStudy (phoenixfire):

According to Wikipedia and several other sites Zero (0) is in fact a multiple of everything.

OpenStudy (jiteshmeghwal9):

so m i correct @PhoenixFire ???

OpenStudy (phoenixfire):

I would say the statement is true. Including zero.

OpenStudy (unklerhaukus):

but that definition is inconsistent with other definitions, that would imply that the lowest common multiple of any numbers would be zero

OpenStudy (jiteshmeghwal9):

yes i'm also with ur opinion.

OpenStudy (phoenixfire):

@UnkleRhaukus No, because the LCM is defined as a positive integer.

OpenStudy (unklerhaukus):

so what is LCM of 10 and 0

OpenStudy (phoenixfire):

If either of the two integers are 0 then the LCM 0.

OpenStudy (phoenixfire):

From Wikipedia: "usually denoted by LCM(a, b), is the smallest positive integer that is divisible by both a and b.[1] If either a or b is 0, LCM(a, b) is defined to be zero."

OpenStudy (unklerhaukus):

well that is clearly a contradiction ,

OpenStudy (phoenixfire):

Hmmm... I don't think that's right. I think the LCM of 10 and 0 would be 10. Majority of the definitions I find agree with one thing: the LCM cannot be 0. However, 0 is still a multiple of everything and is included in the explanation of finding LCMs.

OpenStudy (unklerhaukus):

@Zarkon

Parth (parthkohli):

\[\rm \large (2n + 1)^2 =4n^2 + 4n + 1 = 4(n^2 +n ) + 1 \]

OpenStudy (unklerhaukus):

even?

Parth (parthkohli):

Dude, \(\rm \large n \) can be even.

Parth (parthkohli):

\[\large \rm 4(4k^2+2k)+1=8(2k^2+k)+1\]Proved if n=2k=even

OpenStudy (lgbasallote):

^what did you prove?

Parth (parthkohli):

If n = 2k + 1 = odd, \[\rm \large 4(4k^2+4k+1+2k+1)+1=4(4k^2+6k+2)+1=8(2k^2+3k+1)+1\]

OpenStudy (lgbasallote):

and yes @UnkleRhaukus 0 is a multiple of 8

Parth (parthkohli):

Multiple of 8 + 1 ^^

OpenStudy (unklerhaukus):

im not convinced ,

Parth (parthkohli):

Why not?

OpenStudy (lgbasallote):

mathematicians never are

OpenStudy (lgbasallote):

unless things become too complicated to understand....only then do they agree (when they can't understand it anymore)

OpenStudy (lgbasallote):

i think it's a mathematician's psychological disorder

Parth (parthkohli):

The square of any odd number is given by (2n + 1)^2 where n can be of any parity. = 4n^2 + 4n + 1 Because n can be of any parity, we'd first prove it for odd, then even which I did above.

OpenStudy (lgbasallote):

masochism

OpenStudy (unklerhaukus):

because im not convinced i know the definition of '' is a multiple of '' that is intended in the question

OpenStudy (lgbasallote):

but i do agree with @UnkleRhaukus though... @ParthKohli 's proof is incomplete

Parth (parthkohli):

How so? Do you want one full post of this proof?

OpenStudy (lgbasallote):

1) that's impossible 2) refer to the definition

OpenStudy (lgbasallote):

you'll have to let 2k^2 + 3k = x and then rewrite 8x + 1 <--only then will that proof be complete

OpenStudy (lgbasallote):

it's a formality in proving...you won't understand....

Parth (parthkohli):

The square of any odd number is given by (2n + 1)^2 where n can be of any parity. = 4n^2 + 4n + 1 = 4(n^2 + n) + 1 Because n can be of any parity, we'd first prove it for odd, then even. Case 1 - EVEN: if n is even, then n = 2k where k is another integer. We have 4(n^2 + n) + 1 sub n = 2k, 4(4k^2 + 2k) + 1 = 8(2k^2 + k) + 1 That's a multiple of 8 + 1. Case 2 - ODD: if n is odd, then n = 2k + 1 where k is another integer. We have 4(n^2 + n) + 1 = 4((2k + 1)^2 + 2k + 1) + 1 = 4(4k^2 + 4k + 1 + 2k + 1) + 1 = 4(4k^2 + 6k + 2) + 1 factoring 2 out = 8(2k^2 + 3k + 1) + 1 Multiple of 8 + 1 ^^

Parth (parthkohli):

QED

Parth (parthkohli):

@lgbasallote I've manipulated and rewritten it as 8x + 1 in both cases.

Parth (parthkohli):

@UnkleRhaukus

OpenStudy (lgbasallote):

you still haven't

Parth (parthkohli):

OK. EVEN: let x = 2k^2 + k then we are left with 8x + 1 ODD: let x = 2k^2 + 3k + 1 then we are left with 8x + 1

Parth (parthkohli):

Wow, this is the first time I successfully proved anything by myself.

OpenStudy (unklerhaukus):

i still believe false

Parth (parthkohli):

What else?

OpenStudy (anonymous):

ParthKohli's proof is correct.

OpenStudy (unklerhaukus):

i am not convinced that can zero be considered as a multiple of eight

Parth (parthkohli):

lol

Parth (parthkohli):

All multiples of 8 are in form 8k where k is an integer. 8k = 0 k = 0 and k is an integer QED #2

OpenStudy (anonymous):

Here's an interesting fact : 1^2 = 8*0 +1 3^2 = 8*1 +1 5^2 = 8*3 +1 7^2 = 8*6 +1 9^2 = 8*10 +1 11^2 = 8*15 +1 ... 1,3,6,10... are triangular numbers of the form (1/2)n(n+1), but zero is not a triangular number . the triangle numbers can be related to the square numbers by \[(2n+1)^{2}=8T _{n}+1\]= \[T _{n-1}+6T _{n}+T _{n+1}\] \[T _{n}\] is a triangular number